OpenStudy (meilendurcer):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). A. y= 1over3x + 7 B. y= -3x + 3 C. y= 1over3x + 3 D. y= -3x + 7 Please help!!!

OpenStudy (anonymous):

please give me a medal its B. @meilendurcer

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Okay, first things first. In order for a line to be parallel to one another, they must both have the SAME SLOPE. In this case , y=-3x+12, the slope is 3. So to find the equation of the other line whih passes through the points (-1, 6 ), substitue these into this , y=mx+c. ( Remember the equation for any straight -line graphs are in the form of y=mx+c.) Because ur line has to passthrough the points (-1, 6 ) you have to substitute these points in: so now it becomes 6= -3(-1)+ c ... -3 is your slope, 6 is your y, and -1 is your x. Then just solve it. You'll get c=3. Now you know what all your variables are, aand can write your equation for the seond line now. Y=-3x+3. I just plugged everything in y=mx+c form . Remember, m is the slope and c is the y-intercept. I hope this helped and I didn't just confuse you.